Circular No. 7333

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gl IN NGC 317B G. M. Hurst, Basingstoke, England, reports the discovery by Tom

Boles, Wellingborough, of a supernova (mag 16.2) on unfiltered CCD images obtained in the course of the U.K. Nova/Supernova Patrol on Dec. 14.73 and 14.98 UT with a 0.36-m reflector. Boles measured the position of SN 1999gl to be α = 0^{h}57^{m}40^{s}.07, δ = +43^{o}47′35′′.6 (equinox 2000.0; mean of two exposures). SN 1999gl does not appear on an image obtained by Boles on Sept. 17.02 (limiting mag ∼ 18). M. Armstrong, Rolvenden, confirmed the presence of SN 1999gl on Dec. 14.976 and notes that it was not present on his own CCD image taken on Dec. 4. The second Palomar Sky Survey does not show anything at the position of SN 1999gl (limiting mag ∼ 20). Boles found the supernova to be at mag 16.6 on Dec. 15.92.

P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999gl was obtained by P. Berlind with the Fred L. Whipple Observatory 1.5-m Tillinghast telescope on Dec. 16.1 UT, showing the supernova to be a type-II event near maximum. Broad Hα and Hβ emissions, as well as He I 587.50nm, are evident in the spectrum. A deep Na I absorption line due to interstellar gas in the host galaxy has an equivalent width of 0.3 nm, implying that the supernova is heavily extinguished by dust. Strong, narrow emission lines from the host galaxy provide a redshift of 0.018.

GRB 991208 M. Bremer, Institut de Radio Astronomie Millimétrique (IRAM), Gre-

noble; F. Bertoldi, Max-Planck-Institut für Radioastronomie, Bonn; U. Lisenfeld, IRAM, Granada; Alberto Castro-Tirado, Laboratorio de Astrofísica Espacial y Física Fundamental del Instituto Nacional de Técnica Aerospacial Madrid, and Instituto de Astrofísica de Andalucía, Granada; and T. Galama, California Institute of Technology, report: “The afterglow of GRB 991208 was detected at 240 GHz with the Max-Planck Millimeter Bolometer (MAMBO) array (80-GHz bandwidth) at the IRAM 30-m telescope on Pico Veleta, at the position reported by Frail et al. (IAUC 7332). Three on-off integrations yield the following flux densities: Dec. 11.77 UT, 2.6 ± 0.8 mJy; 12.38, 1.9 ± 0.6; 14.35, 2.2 ± 0.7. Combining all data, we obtain a flux density 2.4 ± 0.5 mJy. The source intensity appears to remain constant. We will continue to monitor its flux.”

1999 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green